OpenStudy (anonymous):

Use the rational roots theorem to form a list of possible rational roots of 3x^7+5x^3-4=0 I have started this problem but am stuck...

OpenStudy (anonymous):

make all possible fractions where the numerator divides 4 and the denominator divides 3

OpenStudy (anonymous):

they look like this \[\pm1,\pm2,\pm4,\pm\frac{1}{3},\pm\frac{2}{3},\pm\frac{4}{3}\]

OpenStudy (anonymous):

hmm ok I was close I had the right idea. thanks how did you decide to use those numbers?

OpenStudy (anonymous):

grind it til you find it start with the easy ones like 1 and -1 btw no one says the actual roots have to be any of these just the POSSIBLE RATIONAL ROOTS

OpenStudy (anonymous):

oh ok I see what you did, makes much more sense then my book! Thank you!